wiki:UC20201006

Version 1 (modified by trac, 16 months ago) ( diff )

--

APO 3.5-m Users Committee Telecon, 10/06/2020

Attending: Nancy Chanover (NMSU), Jamey Eriksen (APO), Bill Ketzeback (APO), John Bally (CU), Moire Prescott (NMSU), Eric Bellm (UW), Kevin Schlaufman (JHU), Anne Verbiscer (UVa), Ben Williams (UW), Misty Bentz (GSU), Denise Stephens (BYU), Russet McMillan (APO), Mukremin Kilic (OU), Joanne Hughes (Seattle U), Sarah Tuttle (UW), Chip Kobulnicky (UWy), Derek Buzasi (FGCU)


User feedback and comments from institutional representatives:

  • FGCU - happy with the nights they had recently
  • NAPG - nothing to report (sent by email)
  • Seattle - thanks for the OPEN time that Joanne was allocated
  • Wyoming - nothing to report
  • Oklahoma - had an issue where a user couldn't use DIS one night because the contamination on the red side was really bad after coming off the vacuum pump; this is discussed below
  • Georgia State - nothing to report
  • UVa - one user requested to have password changes less frequently because quarterly changes seems excessive. Russet replied that this was implemented for security purposes and is necessary. However, if users just need the password in order to download their data, there is a read-only FTP directory from which they can download their data. Russet will get the name of the user from Anne and explain it. All UC reps should remind their users that this option exists. If users have questions they can contact the observing specialists. ARCSAT users still need to use the quarterly institutional passwords.
  • Colorado - thanks to the 3.5m staff for helping out with John Bally's recent student training observing run on ARCSAT
  • JHU - nothing to report
  • NMSU - nothing to report
  • Washington - excited about some more UW people getting trained (taking place this week!)
  • BYU - looking forward to their training run in November

Telescope and Instruments Report:

Jamey's detailed report is included below, followed by additional information discussed during today's call.

3.5-m Telescope and Instruments Highlights, 8/28/2020 - 10/01/2020

0) Overview

APO has been open for night operations since June 1. We are still restricting all visitors to the site. Following the state of New Mexico and NMSU guidelines, there are travel restrictions for staff as well as visitors. Currently, as NMSU is in Phase 3 of the Return to Research, APO is allowed between 35-50% of staffing. The NMSU numbers, coupled with the state of New Mexico restrictions for a non-essential classified business like APO, which allow staffing levels of 50% of the Fire Marshall determined maximum occupancy of a building, has allowed us to have more people on site at any given time. However, those restrictions are still limiting work activities for daytime engineering tasks. Weather for the month of September has been as expected for an average year. We are experiencing an above average year of Miller moths and as such are still dealing with them. The moth issue that abated considerably during monsoon rains has started to get worse again and will continue through the first frost when the moths will crawl into inconvenient places and shelter inside the warm buildings.

1) Telescope

The telescope has been performing as expected. The control system for tertiary tilt had an issue recently. As different ports have different tertiary tilt positions, observing that night was limited to the echelle port. Problem was resolved the next day with a power cycle of the tertiary mirror controller.

2) Instruments

DIS: Scattered light is still an issue with the instrument. After being pumped, there was some alarm about the red camera, but it quickly improved and is now in a very good state with minimal scattered light. The blue camera is being vacuum cycled now and we should have an update by the time of the UC meeting. History tells us that blue may deteriorate once back in service, the question is how rapidly.

TripleSpec: powered on, cold and operational. The pattern noise noted last year is still present at very low level illumination when the instrument is upside down. If this is a matter of concern, ask the observing specialist for tips on how to keep the instrument upright while observing.

Agile: powered on, cold and operational.

Echelle: powered on, cold and operational. The inter-order light (IOL) for red and blue continues to improve slowly. Next instrument service is planned for spring.

NIC-FPS: Frequent expose inits are being required to remedy an all-zero readout problem that occurs more frequently during instrument mounting or movement and then taper off after instrument has been in use for a few hours. Investigations are ongoing.

ARCTIC: powered on, cold, and operational. Issue with dome flats is being debugged and solutions are being worked on. (“Fence” baffle on the mirror cover petal was tested recently and the results are promising.)

3) ARCSAT

ARCSAT is operational and has been used for observations this past month. Driver/software issues associated with changing CCD set point on the TEC cooler as well as upgraded control software have caused some issues. Shelyak eShel commissioning took place at the beginning of September. Data are being analyzed and discussed from the first light work. Further Shelyak Engineering/commissioning will take place at the end of October.


Additional telescope and instrument discussion:

No updates regarding telescope or instruments beyond what was included in the monthly report. The site is still operating at a reduced on-site staffing level due to COVID restrictions. The DIS red problem that the OU user experienced last month resolved itself in that the scattered light improved. The DIS blue camera was just vacuum cycled. It came off the pump looking worse, it improved rapidly, and now is showing signs of getting worse again. Worst case scenario it will return to the state it was in before pumping. As a reminder, several months ago we stated that we would no longer service DIS Blue because after pumping it would revert to major scattered light problems within a matter of hours. With the colder temperatures that the site is now experiencing we decided to try again, but with relatively little hope that it would be greatly improved. In summary, the DIS status is roughly the same as it has been for the last 3 months or so. The DIS Red camera is serviced about once a year so we should not have to do it again for a while.


APO and the COVID-19 virus

The New Mexico Governor's public health order still mandates a 14-day self-quarantine for anyone traveling into New Mexico from out of state. Of relevance to our users, the main issues are (a) we are still not allowing non-essential visitors to the site due to the 14-day quarantine, and (b) we are still operating at a reduced staffing level and with additional cleaning protocols, so please continue to exercise patience with the observing staff. For essential visitors who must come to the site, there is a testing exemption where if one is tested within 72 hours of traveling to NM from out of state (or arriving in NM from out of state) and that test is negative, one does not need to quarantine. We have not investigated on-site or near-site testing options because so far there has not been a need.


Q4 3.5m scheduling:

The Q4 schedule is posted. There is some OPEN time available, particularly B halves. UC reps: notify users that there is a limited amount of OPEN time Q4. If interested in requesting some time, please follow standard protocols for making such requests.


Q4 ARCSAT scheduling:

There was a high demand for ARCSAT in Q4. All weeks have been assigned to requested programs or engineering, thus there is no open time remaining.


Updated Training Policies and Procedures:

We are now half-way through the Q4 pick-up training session (the pre-observing sessions have taken place, and the actual observing will take place later this week). We envisioned these once-per-quarter pickup training sessions to be geared towards new hires (faculty, postdocs, research scientists) or very advanced students, but some faculty members expected their first-year graduate students to be eligible to participate. In the event that students are going to be encouraged to sign up for these trainings, their advisors must be heavily involved. We will report at a future meeting what we expect that involvement to look like. In any case, we strongly encourage faculty to continue to use class groups to get students trained since we can target those trainings to be at a more appropriate level. Nancy took the action item to develop a survey form to get feedback on the virtual training sessions.


KOSMOS Update:

The TEG has had ongoing issues with new larger KOSMOS dewar not holding vacuum. However, they just a ran successful test where the dewar sustained a good vacuum while cold, which is reason for cautious optimism. The next step is to rebuild the innards (minus the detector) and repeat the test. The hope is that within the next 4 weeks they will have the camera fully assembled to do optical testing in lab. They are also prepping the smaller dewar as a backup plan. It may be possible to have it on the mountain for Q1 2021. Sarah will notify us if that changes. The commissioning will require coordination with the mountain, and a commissioning plan is being developed. We may not schedule commissioning time in Q1 unless it looks extremely likely, but we can use engineering time if need be.


ARCTIC Diffuser Rotation

The diffuser rotation mechanism is just about ready for shipment to APO for commissioning.


DAWG Update

Nancy received the final report from the DAWG and will send it out to this group right after the call. It is also accessible from the DAWG page on this wiki. It contains several recommendations ranging from data storage to the migration of TUI to python3; these will be used to evaluate priorities when it comes to allocation of resources.


Science Communications Update

Nancy and Emily continue to post short turnaround snippets on social media that describe how the telescope was used on a given night. They will contact the program PIs ahead of time to ask for permission and solicit any sample images, spectra, guider images, screenshots of the observing team, or anything else that the users wish to provide. The PIs may also elect NOT to participate in this, or can dictate the amount of information that is shared with the public (e.g. no target names). Nancy and Emily also welcome unsolicited submissions for this effort! UC reps: please give your users a heads-up that they may see some of these requests in the near future. Also, a Teen Science Cafe outreach event, where kids would get to see ARCSAT in action, was held last month. The weather was not bad enough to activate the rain date, but not good enough to get nice observations. We will be attempting this again in November in conjunction with an APOLLO run. If anyone is interested, these events are recorded and available for viewing on the Las Cruces Museums Facebook page.


Web Site Tiger Team

A small group is developing a plan for updating the APO web site. They will be identifying representatives from all stakeholder groups to help with this. While they are not seeking general volunteers, if any individuals who would be especially well suited for this project can be identified please let Ben know. Shane has developed a sample site to build out. This is a complex problem with internal apps being developed as well as a desired change to the front-end look and feel. Several UC members requested that every effort be made to make the transition as seamless as possible. If some content will be password protected while other content will not be, we need to make sure that everything is still accessible to the relevant user groups.


ARC Board of Governors Meeting

Nancy reminded the group that the ARC BoG meeting is November 2. So far she has received NO science highlights from anyone in the ARC user community. After some discussion about the purpose of the science highlights, the call was broadened to include feedback from each institution about ANY aspect of ARC membership that they found to be beneficial over the past year. This could include science highlights but it could also be the training of students, the ability to accommodate guest instruments, etc. ANY of those things could be highlighted in a slide, but SOMETHING should be highlighted from each institution! UC reps should remind their users that all slides are due to Nancy by October 15.


ACTION ITEMS:

Open action items from previous meetings:

  • ARCTIC filter order - Nancy. Status: OPEN. She is trying to reconcile some vastly different price quotes from several vendors and is looking to place the order within the next month.
  • UC reps: Remind users that Nancy needs science highlights from each institution for the ARC Board of Governors meeting (see above for broader scope of slides). They are due to Nancy no later than October 15, and can be collected by a single representative from each institution and mailed as a packet (easier for Nancy!) or sent to her directly by individual investigators. There is no slide template, but each slide should contain at minimum: project title, author(s) and affiliation(s), a graphic, a few text bullets describing the result, and a citation if there is one. Status: OPEN.
  • UC reps: alert your users to the fact that they may be asked for contributions to post on social media. Status: OPEN.

New action items from this meeting:

  • UC reps: remind your users that if they need to download their data there is a read-only FTP site from which they can do so without needing the institutional password for the current quarter.
  • UC reps: notify users that there is a limited amount of OPEN time Q4. If interested in requesting some time, please follow standard protocols for making such requests.
  • Nancy: develop a survey to get feedback on the pickup training session.

All Other Business

Nancy raised the issue of whether or not we should participate as an exhibitor in the Jan. AAS meeting, which will be virtual. Some were not excited about participating in a virtual meeting. Some thought that students would be more likely to attend than faculty so maybe we could organize a meet-and-greet for student users and recently trained 3.5m users so they could share their experiences. Nancy is unclear what we could present as a webinar if given the option. She will take this feedback to the Executive Committee; a decision needs to be made by October 22.


Next meeting:

The next meeting will be on Tuesday November 10 at 10:30 am MST.


Note: See TracWiki for help on using the wiki.